# Advection of the image point in probabilistically-reconstructed phase spaces

### v2

@@ 2 The probabilistic reconstruction method: schematic and basics
where $d$ is the Euclidean distance between the $i$-th point in $A$ and a point in $B$; ||$\cdot$|| ⋅ | is the number of points in the cloud (it is assumed that $|A|=|B|$). The cloud A represents the reference data, while the cloud B is the HP solution computed in the probabilistically-reconstructed phase space; the clouds are not supposed to have any order defined on them.
@@
(f) Evolution in physical space. The fifth step is to compute the modelled solution back in the physical space (Fig. 1f). In this study, the solution computed with the HP method is reshaped from a vector (used in the phase space) into a scalar field, which is used to present fields in the physical space.
@@ 3.1 Chua’s circuit
with $\mathbf{x}(t)=(x(t),y(t),z(t))$, and $\alpha=10$, $\beta=15$, $\gamma=0.01$, $a=0.1$, $c=-0.2$. As an initial condition, we take $\mathbf{x}(t_{0})=(1,0,0)$ so that the solution is close to the attractor. Note that the solution $\mathbf{x}(t)$ at time $t$ (being voltages across the capacitors and the current through the coil) is in the physical space, while the whole evolution of the solution (together with the time derivatives of the solution) is in the phase space.
@@
Then, the tendencies are computed by differentiating $\mathbf{x}(t)$; we use the central finite difference in time (Fig. 4). Instead of differentiating the solution, one can use the right hand side, but we recommend against doing that, as the right hand side is usually unavailable for numerical simulations and observations. Note that instead of or in combination with the numerical solution, one can also use observations. In this case, one should form a new dataset $\widehat{\mathbf{x}}=\{\mathbf{x}(t),\mathbf{x}_{o}(t)\}$, which includes both the solution $\mathbf{x}(t)$ and observations $\mathbf{x}_{o}(t)$.
Calculation of JPD and sampling from JPD. The JPD of the reference solution $\mathbf{x}(t)$ and reference tendencies $\mathbf{F}(\mathbf{x}(t))$ are calculated as described above. In order to make sure that sampling from the JPD is accurate enough for further use, we compare the reference PDF (computed from the reference dataset) and the reconstructed PDF (computed from the data sampled from the JPD). As seen in Fig. 5, the reconstructed PDF (blue) is an accurate approximation to the reference PDF (black); the reconstructed PDF can be calculated more accurately, but it is not required for the purpose of this study.
Figure 5: Shown is the coordinate-wise reference PDF (black) and reconstructed PDF (blue) for the solution of Chua’s system (top row) and for the corresponding vector field (bottom row). Note that the reconstructed PDF is an accurate approximation to the reference PDF.
The next step is to sample from the JPD to probabilistically reconstruct the reference phase space for the HP method. We use the same amount of reference data as before, but harvest twice as much from the JPD (Figs. 6,7).
@@
Figure 7: The same as in Fig. 6, but for the tendencies. The tendencies are not sampled directly from the JPD but are computed locally in the neighbourhood of the sampled solution.
The results clearly demonstrate the high quality of the probabilistically-reconstructed data, and that it covers a larger phase space compared with the reference one.
@@ 3.2 Calibration of hyper-parameters
The calibration procedure consists of minimizing $\mathrm{Di}(A,B)$ in the space of hyper-parameters $M$, $N$, and $\eta$. One can opt for engaging a full-blown optimization to find the optimum solution. However, for the purpose of this study, we set $M=N=5$ and search for $\eta\in[0,1]$ that minimizes $\mathrm{Di}(A,B)$. We have found that $\eta=0.235$ gives a solution with low discrepancy.
In order to demonstrate that the proposed method can gain from using the probabilistically-reconstructed phase space, we compare it with the HP method run on the reference dataset (Fig. 8).
@@
As seen in Fig. 8b, the HP solution computed in the probabilistically-reconstructed phase space is very-well developed on the attractor compared with the HP solution computed on the reference dataset (Fig. 8c), which remains bounded to a very narrow band because of lack of data. All these results give an extra layer of confidence that the proposed method has strong potential for the use in more sophisticated models which we consider in the next section.
3.3 Corrupted reference data
@@
Corrupted reference data is ubiquitous in the realm of data-driven methods. Reference data sets lacking states in the phase space or having vast voids (regions of no data) can significantly influence the accuracy of data-driven methods or make them inapplicable. Therefore, it is instructive to study what happens when the reference data is damaged. As an example, we consider three different cases of the reference phase space corrupted with: (1) randomly-missed states, (2) gaps, and (3) cuts (Fig. 9). In all cases, sampling from the JPD reconstructs the phase space up to the extent, which is enough for the HP method to work. In the first case, sampling from the JPD densifies the attractor well (top subplot in Fig. 9b). In the second and third cases, however, there are still some voids in the reconstructed phase space (middle and bottom subplots in Fig. 9b) which cannot be completely filled in by sampling from the JPD, but the HP method does reproduce the dynamics on the attractor (similar to the reference one) nonetheless.
@@
Figure 9: Shown are (a) reference data sets with different level of damage (top – (1) 50% of randomly-chosen states has been removed from the original reference data shown in Fig. 8a, middle – (2) original reference data with gaps, bottom – (3) original reference data with three cuts); (b) phase spaces probabilistically-reconstructed from the corresponding corrupted reference data sets; (c) HP solutions computed in the probabilistically-reconstructed phase spaces. In all cases, the HP method can reproduce the solution on the attractor even when sampling from the JPD cannot completely repair the damage. Note that for the HP method we use $\eta=0.003$ for cases (1) and (2), and $\eta=0.051$ for case (3).
@@
The ability of the HP method to work in corrupted phase spaces can be exploited as a fast reanalysis method in which the complex dynamics of a comprehensive ocean model is replaced with the HP solution. The method can be thought of as a dynamic interpolation which can stitch gaps in observational data with the HP dynamics.
@@ 3.4 NEMO model
The output of the global 1/4 (ORCA025-N4001) and 1/12 (ORCA0083-N006) resolution NEMO model is used in this study. The reference solution is the 5-day mean sea surface temperature (SST) and surface relative vorticity (SRV) in the North Atlantic region $[-83^{\circ}W,-20^{\circ}W]\times[27^{\circ}N,55^{\circ}N]$ over the period of 1979-1993; the period of 1994-2007 is used for validation to check how the HP method performs. The reference solution is interpolated from the 1/12-grid onto the 1/4 grid. The lower-resolution 1/4-solution from ORCA025-N4001 simulation (we call it the modelled solution) is used for comparison with the HP solution.
In order to calibrate the hyper-parameters we take $M=N=5$ and minimize $\mathrm{Di}(A,B)$ with respect to $\eta$. We have found that $\eta=0.01$ gives a low discrepancy. After setting up the hyper-parameters, we compute a probabilistically-reconstructed phase space by sampling twice as much compared to the reference record length (Fig. 10).
@@
The dimensionality of the probabilistically-reconstructed space is $n\approx 30000$. One can run the HP method in the high-dimensional phase space, but in this work we further reduce it down to $m=200$ by using the EOF-PC decomposition [28]. It results in a two-order of magnitude reduction compared with the original dimension; note that the 200 leading EOFs capture 99% of the total SST variance and 60% of the total SRV variance. We compute the EOF-PC decomposition of SST and SRV separately. However, it can be computed for both if one needs to model these fields together.
The space of PCs (let it be denoted as $\mathbf{x}$) can be regarded as a reduced version of the probabilistically-reconstructed phase space. Given this reduced space, one can reconstruct a dynamical system governing the reduced dynamics of observed reference data, as we did in [23]. In this study, we use a different idea: the dynamics of PCs is modelled in the reduced space with the HP method directly and then the HP solution is projected back into the full-dimensional probabilistically-reconstructed phase space.
@@
$\frac{d\mathbf{z}}{dt}=\frac{1}{M}\sum\limits_{j{\color[rgb]{0,0,0}\in}\mathcal{U}_{J}}\mathbf{F}(\mathbf{x}_{j})+\eta\left(\frac{1}{N}\sum\limits_{i{\color[rgb]{0,0,0}\in}\mathcal{U}_{I}}\mathbf{x}_{i}-\mathbf{z}(t)\right),\quad\mathbf{z}(t_{0})=\mathbf{z}_{0},$ (4)
@@
We report the results in Fig. 11 which clearly shows that the HP solution (Fig. 11b) much better represents the Gulf Stream than the modelled solution (Fig. 11c). These results are also confirmed in Fig. 12. Namely, the difference between the time mean of the reference solution and HP solution is much smaller (Fig. 12a) than that of the modelled solution (Fig. 12b). The root mean square error for the HP solution (Fig. 12c) is also much lower then the one of the modelled solution (Fig. 12d). The time mean energy spectral density of the HP solution is closer to the reference than that of the modelled solution over the whole range of wave numbers (Fig. 12e, compare the blue and magenta graphs with the black one), and this difference is even more pronounced for SRV (Fig. 14e).
@@
Figure 12: Shown are (a) difference between the time mean reference SST and HP solution, (b) difference between the time mean reference and modelled SST, (c) root mean square error between the reference SST and HP solution, (d) root mean square error between the reference and modelled SST, (e) the natural logarithm of the time mean energy spectral density $\langle ESD\rangle$ vs the magnitude of the wave vector $\mathbf{k}$. We use $m=200$ in the EOF-PC decomposition for (a)-(d). The $\langle ESD\rangle$ of the HP solution (red and magenta) is closer to the reference (black) than that of the modelled solution (blue). Note that $\langle ESD\rangle$ for $m=200$ (red) and $m=400$ (blue) are almost the same, as adding more EOFs does not contribute much to the accuracy of SST. For illustrative purposes, we apply a locally estimated scatterplot smoothing (also known as a Savitzky–Golay filter) to smooth the ESD [29].
@@
More insights about how the HP solution differs from the modelled solution can be seen in the SRV fields (Fig. 13). The Gulf Stream computed with the HP method (Fig. 13b) is much stronger than the one computed with the 1/4 model (Fig. 13c). Moreover, the HP solution is teemed with vortices (like the reference one), while the vortex dynamics is significantly inhibited in the modelled solution as it is also reflected by the error plots in Fig. 14, as well as by the time mean energy spectral density, denoted as $\langle ESD\rangle$. The latter is more sensitive to the number of EOFs compared with SST, i.e. the more EOFs are used, the more accurate the $\langle ESD\rangle$ is. It is because SRV fields contain much more small-scale features (compared with SST), which therefore require more EOFs to be accurately represented. Doubling the number of PCs ($m=400$) results in a more pronounced Gulf Stream dynamics and larger population of vortices (Fig. 13), as well as in a more accurate representation of $\langle ESD\rangle$ (Fig. 14e). However, it is contribution to the time mean flow and the root mean square error is small (not shown).
Figure 13: The same as in Fig. 12 but for SRV ($m=200$ and $m=400$).

### v3

@@ 2 The probabilistic reconstruction method: schematic and basics
Another important aspect of the HP method is that only first moments of the local reference neighborhood are used: the mean position $\overline{\mathbf{x}}$ and the mean tendency $\overline{\mathbf{F}}(\mathbf{x})$. This yields a trajectory driven by the average local flow and attracted towards the centroid of nearby states, thus offering a stable, low-variance approximation. However, this formulation is agnostic to the geometry of the neighborhood, as it treats all directions equally and fails to exploit any local anisotropy or alignment in the data. Including second-moment information, such as the covariance matrix of positions $\mathbf{C}_{\mathbf{x}}$ and the covariance of tendencies $\mathbf{C}_{\mathbf{F}}$ introduces important geometric structure. For example, the nudging term can be replaced with
$\eta\mathbf{C}^{-1}_{\mathbf{x}}\left(\frac{1}{N}\sum\limits_{i\in\mathcal{U}_{I}}\mathbf{x}_{i}-\mathbf{y}(t)\right)$
to modulate correction strength based on local variance, and the mean tendency can be refined by projecting $\mathbf{F}$ along the principal directions of $\mathbf{C}_{\mathbf{F}}$, thus leading to directionally aware transport, i.e. rather than treating all directions equally, the method adapts to the local flow geometry by steering the trajectory along the locally dominant dynamical pathway. This enhances reconstruction fidelity in phase-space regions where the reference neighborhood is locally stretched, sheared, or curved, which is a common situation in GFD. The downside is that second moments are more susceptible to sampling noise, can be ill-conditioned and require matrix inversion, thus increasing computational cost and risking numerical instability. Furthermore, if the neighborhood is undersampled or nearly isotropic, second-moment corrections may amplify noise without improving accuracy. In short, while the first-moment HP equation is robust, efficient, and broadly applicable, the second-moment extension provides anisotropic, geometry-aware refinement, which is beneficial when the local structure is well-resolved and physically meaningful, but risky when data is sparse or noisy.
@@
where $d$ is the Euclidean distance between the $i$-th point in $A$ and a point in $B$; $|\cdot|$ is the number of points in the cloud (it is assumed that $|A|=|B|$); $\|\cdot\|_{2}$ represents the $l_{2}$-norm. The cloud A represents the reference data, while the cloud B is the HP solution computed in the probabilistically-reconstructed phase space; the clouds are not supposed to have any order defined on them.
@@
(f) Evolution in physical space. The fifth step is to compute the modelled solution back in the physical space (figure 1f). In this study, the solution computed with the HP method is reshaped from a vector (used in the phase space) into a scalar field, which is used to present fields in the physical space.
@@ 3.1 Chua’s circuit
with $\mathbf{x}(t)=(x(t),y(t),z(t))$, and $\alpha=10$, $\beta=15$, $\gamma=0.01$, $a=0.1$, c=0.2c=-0.2. As an initial condition, we take $\mathbf{x}(t_{0})=(1,0,0)$ so that the solution is close to the attractor. Note that the solution $\mathbf{x}(t)$ at time $t$ (being voltages across the capacitors and the current through the coil) is in the physical space, while the whole evolution of the solution (together with the time derivatives of the solution) is in the phase space.
@@
Then, the tendencies are computed by differentiating $\mathbf{x}(t)$; we use the central finite difference in time (figure 4). Although the right-hand side is explicitly known for Chua’s circuit, we do not use it in this study. Instead, we estimate tendencies via central finite differences applied to the numerical solution in order to ensure methodological consistency with realistic applications, where the exact right-hand side may not be available. The HP method is designed to operate solely on state-tendency pairs extracted from data, without relying on knowledge of the underlying equations. Also note that instead of or in combination with the numerical solution, one can also use observations. In this case, one should form a new dataset $\widehat{\mathbf{x}}=\{\mathbf{x}(t),\mathbf{x}_{o}(t)\}$, which includes both the solution $\mathbf{x}(t)$ and observations $\mathbf{x}_{o}(t)$. However, using observations in the HP approach is not just adding observed points into the phase space, unless they lie on the same attractor or in its near vicinity (it is doubtful due to model biases, resolution differences, or unresolved physics), they are noise-free (it is never true), they respect the dynamical constraints of the simulation model (it is rare). Put differently, simply injecting observed states into the phase space is insufficient and potentially destructive.
To prevent mismatches between simulations and observations, and a possible drift of the HP solution towards the observational subspace, an additional nudging term should be introduced in equation (1) to steer trajectories towards observations with a tunable strength; this gives the control over how strongly one ”trusts” the observations versus the numerical simulations. When properly balanced against the existing nudging strength $\eta$, this hybrid formulation yields trajectories that should lie between observed and simulated clouds, enabling dynamics that respects both sources. The method’s behaviour depends critically on the geometry, density, and consistency of these two clouds, and on the choice of hyper-parameters. A weighted-neighbourhood formulation is recommended to balance the influence.
Also, observed tendencies can be included in equation (1). However, one should keep in mind that observed tendencies are often noisy, sparse, or inconsistent with the model. The safest route is to use the tendencies from the numerical simulations, while using observations to influence the trajectory via the observation-based nudging term. But, if one does add observed tendencies they should be prperly weighted and smoothed to mitigate their possibly negative effect.
Calculation of JPD and sampling from JPD. The JPD of the reference solution $\mathbf{x}(t)$ and reference tendencies $\mathbf{F}(\mathbf{x}(t))$ are calculated as described above. To ensure that the sampling from the JPD is sufficiently accurate for further use, we compare the reference PDF (computed from the reference dataset) and the reconstructed PDF (computed from the data sampled from the JPD). As seen in figure 5, the reconstructed PDF (blue) is an accurate approximation to the reference PDF (black). While it is possible to calculate the reconstructed PDF with higher accuracy, doing so is not necessary for the purpose of this study.
Figure 5: Shown is the marginal reference PDF (black) and reconstructed PDF (blue) for the solution of Chua’s system (top row) and for the corresponding vector field (bottom row). Note that the reconstructed PDF is an accurate approximation to the reference PDF.
The next step is to sample from the JPD to probabilistically reconstruct the reference phase space for the HP method. We use the same amount of reference data as before, but harvest twice as much from the JPD (figures 6,7).
@@
Figure 7: The same as in figure 6, but for the tendencies. The tendencies are not sampled directly from the JPD but are computed locally in the neighbourhood of the sampled solution.
The results clearly demonstrate not only the high quality of the probabilistically-reconstructed data, but also that the reconstructed data covers a larger phase space compared with the reference one. In other words, sampling from the JPD reconstructs the reference space geometry thus implying correct dynamical behaviour of the HP method, i.e. the solution has the right range of variability and local curvature, not just the right mean or energy.
@@ 3.2 Calibration of hyper-parameters
The reconstructed dynamics in phase space, computed with the HP method, depends on the hyper-parameters $M$, NN, $\eta$. Our studies (not shown) allow us to report the following findings. Small values of $M$ and $N$ yield high-variance estimates of local means and tendencies, which in turn can produce noisy, over-dispersed trajectories and inflated attractors. Conversely, large $M$ and $N$ lead to over-smoothed reconstructions, suppressing local variability and causing the attractor to shrink unnaturally. The nudging parameter $\eta$ controls the strength of the restoring term that pulls the trajectory toward the local mean: small $\eta$ allows freer movement through phase space and results in broader, more variable attractors, while large $\eta$ imposes strong corrections that constrain the dynamics, often collapsing the trajectory onto the reference one. The balance of these parameters governs the trade-off between preserving dynamical richness and maintaining numerical stability in the reconstruction.
In order to properly adjust these hyper-parameters, one can calibrate them by minimizing $\mathrm{Di}(A,B)$ in the space of $M$, NN, and $\eta$. One can opt for engaging a full-blown optimization to find the optimum solution. However, for the purpose of this study, we set $M=N=5$ and perform a one-dimensional grid search over $\eta\in[0,1]$, selecting the value that minimizes $\mathrm{Di}(A,B)$. We have found that $\eta=0.235$ gives a solution with the lowest discrepancy. Once M and N are fixed, the solution for a given $\eta$ is not necessarily the one that ensures the global minimum, only a local one. Currently, only the accuracy in reconstructing the reference phase space is considered. But in principle, one could also optimize for temporal smoothness of the HP trajectory, physical consistency, forecast skill, robustness to perturbations, etc.
In order to demonstrate that the proposed method can gain from using the probabilistically-reconstructed phase space, we compare it with the HP method run on the reference dataset (figure 8).
@@
As seen in figure 8b, the HP solution computed in the probabilistically-reconstructed phase space is very-well developed on the attractor compared with the HP solution computed on the reference dataset (figure 8c), which remains bounded to a very narrow band due to insufficient data. All these results give an extra layer of confidence that the proposed method has strong potential for the use in more sophisticated models which we consider in the next section.
3.3 Damaged reference data
@@
Reference data with unreliable or missing states (which are called damaged in this study) is ubiquitous in the realm of data-driven methods. Reference data sets lacking states in the phase space or having vast voids (regions of no data) can significantly influence the accuracy of data-driven methods or make them inapplicable. Therefore, it is instructive to study what happens when the reference data is damaged. As an example, we consider three different cases of the reference phase space damaged with: (1) randomly-missed states, (2) gaps, and (3) cuts (figure 9). In all cases, sampling from the JPD reconstructs the phase space up to the extent, which is enough for the HP method to work. In the first case, sampling from the JPD densifies the attractor well (top subplot in figure 9b), because the global support of the attractor (the bounded region in phase space within which all trajectories of the dynamical system ultimately reside and evolve) remains intact, and only the sampling density is reduced. In the second and third cases, however, there are still some voids in the reconstructed phase space (middle and bottom subplots in figure 9b) which cannot be completely filled in by sampling from the JPD, because the global support of the attractor is significantly distorted by removed coherent regions of the attractor, but the HP method does reproduce the dynamics on the attractor (similar to the reference one) nonetheless.
@@
Figure 9: Shown are (a) reference data sets with different level of damage (top – (1) 50% of randomly-chosen states has been removed from the original reference data shown in figure 8a, middle – (2) original reference data with gaps, bottom – (3) original reference data with three cuts); (b) phase spaces probabilistically-reconstructed from the corresponding damaged reference data sets; (c) HP solutions computed in the probabilistically-reconstructed phase spaces. In all cases, the HP method can reproduce the solution on the attractor even when sampling from the JPD cannot completely repair the damage. Note that for the HP method we use $\eta=0.003$ for cases (1) and (2), and $\eta=0.051$ for case (3).
@@
The ability of the HP method to work in damaged phase spaces can be exploited as a fast reanalysis method in which the complex dynamics of a comprehensive ocean model is replaced with the HP solution. Essentially, this method can be thought of as a dynamic interpolation, which can stitch gaps in observational data using the HP solution.
On the other hand, linear or spline interpolation can fill gaps between known samples too, but these methods operate purely in the static data space and do not account for the underlying dynamics of the system. In contrast, the HP method reconstructs missing states by integrating an explicit dynamical model in the reconstructed phase space, guided by the joint probability structure of the reference data. This ensures that the reconstructed trajectories:
  1. 1.
    Remain within the reference phase space, thus avoiding excursions into dynamically forbidden regions that static interpolation can produce;
  2. 2.
    Respect the local flow geometry, as the HP update uses neighbouring states in phase space to determine both the direction and speed of evolution;
  3. 3.
    Handle sparse or irregular sampling, i.e. cases in which simple interpolation becomes ill-defined or distorted.
In the low-dimensional example in figure 9, spline or linear interpolation could visually connect points, but such reconstructions would not match the true time evolution over longer intervals, particularly in cases with under-sampling or in higher-dimensional settings such as SST or SRV, where the reference space geometry is much more complex.
The HP method produces dynamically consistent trajectories because it does not interpolate states purely in a geometric sense. Instead, the method integrates an explicit evolution equation in phase space, whose right-hand side is constructed from local dynamical information in the reference dataset. Specifically, the HP update uses the tendencies $F(x_{j})$ of neighbouring states $x_{j}$ in phase space, together with a relaxation term towards nearby observed states. These neighbours are selected so that they correspond to states actually visited by the system, ensuring that the reconstructed fields are consistent with the reference flow dynamics. By integrating this data-driven dynamical system, the HP method advances the state along directions observed in the reference dynamics, thereby avoiding excursions into dynamically forbidden regions. In contrast, spline or linear interpolation merely connects points geometrically, without enforcing consistency with the underlying vector field, and can therefore produce trajectories that are not dynamically realisable.
Before we proceed to the application of the method to the NEMO model, it is instructive to clarify a subtle point: how the method, although demonstrated to work with data from an autonomous dynamical system, can be equally applicable to data originating from a non-autonomous ocean model. This is possible because the HP method is fundamentally time-agnostic, as it operates in phase space, which contains no explicit time coordinate. Whether the system is autonomous or non-autonomous is irrelevant to the method’s mechanics. What matters is that the reference dataset contains a representative sample of the system’s possible states, including those generated under time-varying forcings. If so, the HP trajectory can reconstruct dynamical behavior from the geometry of the reference phase space. Time enters only indirectly, through how it shapes the distribution of states, not through any explicit appearance in the formulation. However, if the reference phase space does not contain sufficient coverage of the system’s evolution under time-varying forcings, HP cannot reconstruct this dynamics.
@@ 3.4 NEMO model
The output of the global 1/4 (ORCA025-N4001) and 1/12 (ORCA0083-N006) resolution NEMO model is used in this study. The reference solution is the 5-day mean sea surface temperature (SST) and surface relative vorticity (SRV) in the North Atlantic region $83^{\circ}W,-20^{\circ}W]\times[27^{\circ}N,55^{\circ}N]$ over the period of 1979-1993; the period of 1994-2007 is used for validation to check how the HP method performs. The reference solution is interpolated from the 1/12-grid onto the 1/4 grid. The lower-resolution 1/4-solution from ORCA025-N4001 simulation (we call it the modelled solution) is used for comparison with the HP solution.
In order to calibrate the hyper-parameters we take $M=N=5$ and minimize $\mathrm{Di}(A,B)$ with respect to $\eta\in[0,1]$. We have found that $\eta=0.01$ gives the lowest discrepancy. After setting up the hyper-parameters, we compute a probabilistically-reconstructed phase space by sampling twice as much compared to the reference record length (figure 10).
@@
The dimensionality of the probabilistically-reconstructed space is $n\approx 30000$. One can directly run the HP method in this high-dimensional phase space, or in a space of reduced dimensionality if greater speed gains are wanted. In this work, we reduce the phase space dimension down to $m=200$ by using the EOF-PC decomposition [17]; note that we decompose the full flow field, including its time mean. It results in a two-order of magnitude reduction compared with the original dimension; note that the 200 leading EOFs capture 99% of the total SST variance and 60% of the total SRV variance. We compute the EOF-PC decomposition of SST and SRV separately; a bivariate SST-SRV decomposition can be computed if one needs to model these fields together. As an alternative to the bivariate decomposition, one can decompose these fields separately and combine them in the HP method, as in [22].
The space of PCs (denoted as $\mathbf{x}$) can be regarded as a reduced version of the probabilistically-reconstructed phase space. Given this reduced space, one can reconstruct a dynamical system governing the reduced dynamics of observed reference data, as we did in [21]. In this study, we use a different idea: the dynamics of PCs is directly modelled in the reduced space with the HP method and then the HP solution is projected back into the full-dimensional probabilistically-reconstructed phase space.
@@
$\frac{d\mathbf{z}}{dt}=\frac{1}{M}\sum\limits_{j\in\mathcal{U}_{J}}\mathbf{F}(\mathbf{x}_{j})+\eta\left(\frac{1}{N}\sum\limits_{i\in\mathcal{U}_{I}}\mathbf{x}_{i}-\mathbf{z}(t)\right),\quad\mathbf{z}(t_{0})=\mathbf{z}_{0},$ (4)
@@
We report the results in figure 11, which clearly shows that the HP solution (figure 11b) much better represents the Gulf Stream than the modelled solution (figure 11c). These results are also confirmed in figure 12. Namely, the difference between the time mean of the reference solution and HP solution is much smaller (figure 12a) than that of the modelled solution (figure 12b). The root mean square error for the HP solution (figure 12c) is also much lower then the one of the modelled solution (figure 12d). The time mean energy spectral density of the HP solution is closer to the reference than that of the modelled solution over the whole range of wave numbers (figure 12e, compare the blue and magenta graphs with the black one), and this difference is even more pronounced for SRV (figure 14e).
Note that the results shown in figures 11-14 were obtained using the probabilistically-reconstructed phase space generated by sampling from the estimated JPD. Specifically, the reference dataset (1979-1993) was used to estimate the JPD, from which additional states were sampled to densify the phase space prior to EOF reduction and HP integration.
Remark. The HP solution ’much better” or ”significantly better” represents the Gulf Stream because it reconstructs sharper, more coherent, and spatially accurate SST than the native 1/4-model. This improvement arises not from enhanced physics but from the HP method’s ability to statistically emulate high-resolution dynamics within a low-dimensional phase space. Unlike the modelled solution (which suffers from resolution-induced diffusion, under-resolved mesoscale processes, and smeared fronts) the HP solution inherits and preserves fine-scale structure resolved at 1/4. In this sense, ’much better” or ”significantly better” means that the HP trajectory remains geometrically and statistically consistent with the reference solution that the 1/4-model cannot represent.
@@
Figure 12: Shown are (a) difference between the time mean reference SST, $\langle{\rm ref\,SST}\rangle$, and HP SST, $\langle{\rm HP\,SST}\rangle$,, (b) difference between the time mean reference SST and modelled SST, $\langle{\rm model\,SST}\rangle$, (c) root mean square error, RMSE, between the reference SST and HP SST, (d) root mean square error, RMSE, between the reference and modelled SST, (e) the natural logarithm of the time mean energy spectral density $\langle ESD\rangle$ vs the magnitude of the wave vector $\mathbf{k}$; the units in subplots (a)-(d) and (e) are $^{\circ}{\rm C}]$ and $^{\circ}C\cdot{\rm deg}^{2}]^{2}$, respectively. We use $m=200$ in the EOF-PC decomposition for (a)-(d). The $\langle ESD\rangle$ of the HP solution (red and magenta) is closer to the reference (black) than that of the modelled solution (blue). Note that $\langle ESD\rangle$ for $m=200$ (red) and $m=400$ (blue) are almost the same, as adding more EOFs does not contribute much to the accuracy of the $\langle ESD\rangle$ of SST. For illustrative purposes, we apply a locally estimated scatterplot smoothing (also known as a Savitzky–Golay filter) to smooth the ESD [8].
@@
More insights about how the HP solution differs from the modelled solution can be seen in the SRV fields (figure 13). The Gulf Stream computed with the HP method (figure 13b) is much stronger than the one computed with the 1/4 model (figure 13c). Moreover, the HP solution is teemed with vortices (like the reference one), while the vortex dynamics is significantly inhibited in the modelled solution as it is also reflected by the error plots in figure 14, as well as by the time mean energy spectral density $\langle ESD\rangle$, and computed from the two-dimensional Fourier transformation of the fields on the longitude-latitude grid as follows.
Let $q(\lambda,\phi,t)$ denote either SST or SRV at time $t$, where longitude $\lambda\in[\lambda_{\min},\lambda_{\max}]$ and latitude $\phi\in[\phi_{\min},\phi_{\max}]$ are measured in degrees. We denote by $(k_{1},k_{2})\in\mathbb{Z}^{2}$ the integer Fourier indices associated with the zonal (longitude) and meridional (latitude) directions, respectively. We define the corresponding wave-vector components (in units of deg-1) as $k_{\lambda}=k_{1}/L_{\lambda}$ and $k_{\phi}=k_{2}/L_{\phi}$, where $L_{\lambda}=\lambda_{\max}-\lambda_{\min}$, $L_{\phi}=\phi_{\max}-\phi_{\min}$. The wave-vector magnitude is $|\mathbf{k}|=\sqrt{k_{\lambda}^{2}+k_{\phi}^{2}}$. For a given magnitude $|\mathbf{k}|$, we group Fourier modes into a ring defined by $\mathrm{ring}(|\mathbf{k}|)=\left\{(k_{1},k_{2})\in\mathbb{Z}^{2}:\ \sqrt{k_{\lambda}^{2}+k_{\phi}^{2}}=|\mathbf{k}|\right\}$, i.e. $\mathrm{ring}(|\mathbf{k}|)$ contains all discrete Fourier modes having the same magnitude $|\mathbf{k}|$.
Let $\widehat{q}(k_{1},k_{2},t)$ denote the 2D discrete Fourier transform coefficient of $q(\lambda,\phi,t)$ at mode $(k_{1},k_{2})$. For each snapshot $t$, we define the energy spectral density as
$\mathrm{ESD}(\mathbf{k},t):=\sum_{(k_{1},k_{2})\in\mathrm{ring}(|\mathbf{k}|)}\left|\widehat{q}(k_{1},k_{2},t)\right|^{2}.$
The time mean energy spectral density is then
$\langle\mathrm{ESD}(\mathbf{k})\rangle=\frac{1}{T}\sum_{n=1}^{T}\mathrm{ESD}(\mathbf{k},t_{n}),$
where $T$ is the number of available snapshots $\{t_{n}\}_{n=1}^{T}$ in the period used for averaging.
Note that the small-scale oscillations observed for $|\mathbf{k}|>1$ in figures 12,14,16 represent artifacts of spectral estimation, and are not interpreted as physically meaningful spectral features.
SRV is more sensitive to the number of EOFs compared with SST, i.e. the more EOFs are used, the more accurate the $\langle ESD\rangle$ is. It is because SRV fields contain much more small-scale features (compared with SST), which therefore require more EOFs to be accurately represented. Although doubling the number of PCs ($m=400$) results in a more pronounced Gulf Stream dynamics and larger population of vortices (figure 13), as well as in a more accurate representation of $\langle ESD\rangle$ (figure 14e), its contribution to the time mean flow and the root mean square error is small (not shown).
Figure 13: The same as in figure 11 but for SRV ($m=200$ and $m=400$).